OpenStudy (trojanpoem):

An electric motor which gives electric power = 500 W , to turn a pull whose moment of inertia around the the rotation axis is 2 kg.m^2 with speed of 10 revolutions per second. what's the time required to stop the pull after cutting the electric current. Notice that the moment of friction is half after cutting the electricity

OpenStudy (michele_laino):

the energy \(E\) transfer from electric motor to pulley, is: \[E = \frac{1}{2}I\omega _0^2\] where: \[{\omega _0} = 20\pi \]

OpenStudy (trojanpoem):

w = 10 * 2 pi = 20 p i ( Understood) K.E = 0.5 I w^2 K.E = P = 500t ?

OpenStudy (trojanpoem):

0.5 * 2 * (20pi)^2 = 500 t

OpenStudy (trojanpoem):

t = 7.89 s

OpenStudy (trojanpoem):

I think that doesn't work. I got the time needed to get 500 J of energy

OpenStudy (trojanpoem):

we need to make use of the friction force.

OpenStudy (michele_laino):

after the cutting energy, the resistant moment is half of the electric moment, right?

OpenStudy (michele_laino):

so, the power developed by friction is half of the starting power which is \(500\;W\)

OpenStudy (michele_laino):

so we can write this differential equation: \[\huge \frac{{dL}}{{dt}} = - \frac{{{W_0}}}{2}dt\] where \(W_0=500\), and \(L\) is the work done by the electric motor

OpenStudy (trojanpoem):

Moment of rotating = Moment caused by motor - moment of friction

OpenStudy (trojanpoem):

, the power developed by friction is half of the starting power which is 500W ? Why did you induce that ?

OpenStudy (michele_laino):

when I make a cutting of electrical energy, the moment of external forces, which is acting on the pulley, is only the moment related to the friction forces

OpenStudy (trojanpoem):

Can't grasp it well.

OpenStudy (michele_laino):

If there is no electrical energy, what are the external forces which are acting on the pulley?

OpenStudy (trojanpoem):

The friction only

OpenStudy (michele_laino):

and, according to the text of the exercise, the moment due to such friction forces is "...is half after cutting the electricity..."

OpenStudy (trojanpoem):

Let moment of friction before cutting the electricity : \[\tau _{2 } = 0.5 \tau _{1}\]

OpenStudy (michele_laino):

I think that if the new moment is half of the first one, also the absorbed power of such friction forces is half of the power developed by the electrical motor

OpenStudy (michele_laino):

If I integrate the equation above, I get the requested time \(\tau\), as below: \[\huge \tau = \frac{{I\omega _0^2}}{{{W_0}}}\]

OpenStudy (trojanpoem):

Hmm, still not convinced by last part.. Wasn't moment of pully = moment caused by motor - moment of friction ?

OpenStudy (trojanpoem):

The final answer must be 31.55 s

OpenStudy (trojanpoem):

After cutting the current the moment of pully = - moment of friction

OpenStudy (michele_laino):

the moment is referred to the external forces which are acting on the pulley

OpenStudy (trojanpoem):

Pully moment = moment of caused by motor - external forces moment.

OpenStudy (trojanpoem):

I think I am destroying the physics xD

OpenStudy (michele_laino):

if there is not electrical energy, then moment of the electrical motor is zero

OpenStudy (trojanpoem):

so moment of external forces = moment of pulley

OpenStudy (trojanpoem):

moment of pulley = - moment of external forces

OpenStudy (trojanpoem):

moment of external forces after cutting = 0.5 before cutting moment of pulley = - 0.5 moment of external force before cutting

OpenStudy (michele_laino):

please wait a moment, we have to keep in mind that the friction is acting on the pulley also during the working of the electrical motor, namley not all tthe power developed by the electrical motor goes to accelerate such pulley

OpenStudy (trojanpoem):

That's why I said : Pully moment = moment of caused by motor - external forces moment.

OpenStudy (michele_laino):

I know, nevertheless I meant to the second phase

OpenStudy (michele_laino):

during second phase there is the moment due to friction forces only

OpenStudy (trojanpoem):

I am convinced by this

OpenStudy (trojanpoem):

Now ?

OpenStudy (michele_laino):

with my equation I got 31.55/2=15.79 seconds, so please wait...

OpenStudy (trojanpoem):

Do you think your equation will be fixed ?

OpenStudy (michele_laino):

I have to rewrite another equation. Here is my reasoning: during the working of electric motor, we have the subsequent equation: \[\huge {W_0} = {M_f} \cdot {\omega _0}\] where \(W_0=500\), \M_f\) is the moment due to friction force, and \(\omega_0=10\) Please keep in mind during the working of electric motor, we have a stationary or equilibrium situation, namley the moment applied by the motor is against the friction forces only

OpenStudy (michele_laino):

oops.. \(M_f\) is the moment due to friction forces and developed by the same electric motor

OpenStudy (michele_laino):

namely*

OpenStudy (michele_laino):

After the electric energy cutting, the new value of \(M_f\), is: \[\huge \frac{{{M_f}}}{2}\] according to the text of your exercise

OpenStudy (michele_laino):

so we can write the subsequent differential equation: \[\huge \begin{gathered} \frac{{{M_f}}}{2} = \frac{{{W_0}}}{{2{\omega _0}}} \hfill \\ \hfill \\ \frac{{{W_0}}}{{2{\omega _0}}} = - I\frac{{d\omega }}{{dt}} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

next I integrate such equation, so I get: \[\huge \int_0^\tau {dt} = \frac{{ - 2I{\omega _0}}}{{{W_0}}}\int_{{\omega _0}}^0 {d\omega } \]

OpenStudy (michele_laino):

and finally: \[\huge \tau = \frac{{2I\omega _0^2}}{{{W_0}}}\]

OpenStudy (trojanpoem):

Nice.

OpenStudy (michele_laino):

:)

OpenStudy (trojanpoem):

Remember the ice cube question I asked you about ? You can T final as -60

OpenStudy (michele_laino):

I remember, even if, I think that my equation is right

OpenStudy (trojanpoem):

I told you , you were right, the final answer is indeed -60 , But the question asks further more to explain what does -60 as final temperature means.

OpenStudy (michele_laino):

I think that I have made an error, since I used as specific heat of ice the same value of the specific heat of water.

OpenStudy (michele_laino):

I remember that, please check, if I'm wrong, the specific heat of ice is different from the specific heat of water